### v1

@@ 4.3.7 Bend & Twist in Insect-like Wings
Buckling | Sipos and Várkonyi (2020) | CRT | Study relationship between buckling and on simulated cantilever rods. |
@@ 4.4 Bend & Stretch
The simultaneous bending and stretching emerge in many cable-driven and pneumatic CSM (Fig. 5). Table 4 reports illustrative examples of the Bend & Stretch class.
4.4.1 Bend & Stretch by Cables
Deflection and stiffness estimation are crucial for continuum robots actuated by cables. For instance, robot deflection and stiffness were predicted considering environmental loads with constant axial tendon displacements (Oliver-Butler et al., 2019). Preliminary experiments with an Euler–Bernoulli rod evaluate how tendon stretch and routing path affect the robot stiffness at any point. The analytical results enable stiffness analysis of candidate robot designs with light simulations and suggest the ability to increase robot stiffness by using converging tendon paths. Finally, they extend the study to a non-linear Cosserat rod model with large deflections, handling prescribed tendon displacements, tendon stretch, pretension, and slack. Experimental validation on a prototype shows good agreement. Also, a novel stiffness modulation approach using a Cosserat rod model was proposed for a tendon-driven soft CTR (Xiao et al., 2022). The experimental validation shows an accurate shape estimation. They also demonstrate that the robot stiffness changes with a Nitinol backbone. In particular, the tip compliance can be halved, with potential applications in manipulation tasks.
Concerning workspace estimation, the CRT is also applied to investigate equilibrium positions reachable by the end-effector of a planar tendon-driven CSM with a discrete actuation space (Walid et al., 2022). Experiments on a physical prototype efficiently estimated the robot workspace. Another example predicts the soft arm shape combining the EBRT, which considers the influence of external forces in bending, with the PCC model, which efficiently derives a relationship between cable length and arm curvature (Chen et al., 2022). Experiments on a cable-driven soft arm subject to external forces, or , yield a maximum error in the range .
4.4.2 Bend & Stretch in Soft Pneumatic Actuators
A study compares five models, including the Cosserat rod, a dynamic lumped system, and approaches with constant or variable curvature (Sadati et al., 2017b). The comparison made in simulation and on the STIFF-FLOP soft arm (Cianchetti et al., 2013) demonstrated that the lumped mass model better predicts the dynamics, while constant curvature methods are less accurate but are suitable for design optimization. Additionally, they modeled CSM actuated by braided pneumatic actuators (Sadati et al., 2017a). This specific design causes cross-sectional deformation, which was captured analytically through geometry deformation and CRT. This approach outperformed constant curvature approaches by and variable curvature by . A subsequent work presents a semi-analytical model for the dynamics of multi-segment CSM (Sadati et al., 2017c). The Cosserat rod static and Lagrangian dynamic models are derived and combined with the Ritz-Galerking method to reduce the infinite state space of CSM to a small finite number of states.
Moreover, a static Kirchhoff rod model was proposed for a hybrid CSM characterized by pneumatic actuators with embedded elastic rods, computing its deformations using the minimum energy principle (Sun et al., 2019). In addition, the classical CRT modeled a variable-length pneumatic CSM (Gilbert and Godage, 2019). They optimize the cross-sectional area of the actuators and a parameter relating stiffness and damping in the constitutive law. The model is validated experimentally on dynamic bending deformations obtained by pressurizing the actuators with alternating square pulses. A study characterized the bending behavior of a cantilevered extensile fluidic artificial muscle (Garbulinski and Wereley, 2022). An Euler-Bernoulli rod determined static properties like the bending stiffness, which increases more than four times in the pressure interval psi. Experimental data are applied to assess the dynamic properties, such as natural frequency and damping. Also, it was developed an analytical model of a PneuNet actuator based on the Euler-Bernoulli finite strain rod (Sachin et al., 2022). They describe the actuator behavior in free motion and for tip contacts and accurately predict the actuator deformation and the force characteristics. Another work showcases a parallel extensible soft robot with three pneumatic actuators, each with a silicone body, inner air chamber, winding wiring, and sealing head (Wang et al., 2022c). The dynamic model of a single actuator uses the CRT, and an interactive constraint topological structure solves the constraint force and moment of the parallel robot. A linear mapping between the pressure and the extension force is established for each actuator. Validation measures the static step response of a single actuator and conducts dynamic experiments with the complete soft robot. The model accurately captures the actuator dynamics and the parallel robot constraints. Rod models met traditional robotics models where the EBRT augmented the kinematic model of a pneumatic soft bending finger treated as a planar serial rigid robot parameterized with the Denavit-Hartenberg convention (Flores-Martínez et al., 2022). The experimental validation compares the actuator trajectory versus the kinematic model and the flexion-extension motion trajectory of average human fingers.
Manufacturing uncertainty could be significant in fluidic soft actuators. For instance, Eugster et al. (2022) described the kinematics of a soft arm using a CRT with a nonlinear pressure-dependent constitutive law, using the principle of virtual work and modeled the actuator with strain energy functions. Manufacturing imperfections are considered by scaling the relations of extensional and bending stiffness and by a constant shift of the actuator position . Assuming the pressure dilates the chambers, the model also considers cross-section deformations with a pressure-dependent diameter . The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings. They compare models with constant or pressure-dependent chamber radius, with or without chamber repositioning on the cross-section, and with a linear or Ogden material law. The strain energy for an Ogden material (Odgen et al., 2013) is
Here, is a material parameter, and is the principal axial stretch. Similarly, Alessi et al. (2023a) presented a dynamic Cosserat rod model for a 3D-printed pneumatic CSM. The model considers the pressure transient with a parameterized smooth step. Also, it captures the manufacturing variability of the actuators using strain gains that tune the pressure-strain relation for each actuator
where is the interior actuator area. An extensive experimental validation includes quasi-static and dynamic motion patterns at various actuation frequencies. Finally, an ablation study evaluates the contribution of different model components, uncovering that neglecting gravity and manufacturing uncertainties could degrade performance up to , depending on the motion.
Concerning fiber-reinforced actuators, Berthold et al. (2022) introduces a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs. By converting the inner chamber pressure into an equivalent force, they consider the compression effects of lateral pneumatic chamber walls. The validation compared the results to FEM simulations, obtaining a maximum mean tip position error of . Differently, Hanza and Ghafarirad (2023) introduces an inhomogeneous Cosserat rod model for a CSM with pneumatic fiber-reinforced actuators, using a nonlinear strain-force relationship to account for the effect of radial pressure. The proposed method outperformed the standard CRT by about .
Alongside, Liu et al. (2023) discusses another modelling approach of a pneumatic fiber-reinforced soft actuator with a semicircular cross section. Here, the EBRT is employed to explore the effects of self-gravity and external loads on the actuator configuration, to predict its shape.
Dou et al. (2023) introduce a hybrid CSM with pneumatic chambers and an antagonistic rod-driven compliant mechanism enabling elongation, contraction, bending, and stiffness regulation. They derive a kinematic model by adapting the EBRT and demonstrate the benefit of the compliant mechanism with an experimental validation comprising way-point tracking.
Recently, Lamping and de Payrebrune (2023) presented a study on soft robotic actuators that utilize pneumatic actuation with multiple chambers. The modeling approach uses the EBRT to characterize the deformation, providing a practical method for accurately determining the rod parameters essential for predicting the performance of these actuators, enhancing their design and control.
Subclass | References | Rod Model | Contribution |
Cable | Oliver-Butler et al. (2019) | EBRT/CRT | Stiffness analysis for various tendon stretch and routing paths . |
Chen et al. (2022) | EBRT | Augments the PCC with a rod model to quickly solve the CSM deformation under . | |
Pneumatic | Sadati et al. (2017c) | CRT | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. |
Alessi et al. (2023a) | CRT | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | |
Dou et al. (2023) | EBRT | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | |
Smart | Goldberg et al. (2019) | CRT | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. |
4.4.3 Bend & Stretch with Smart Actuators
Regarding locomotion with smart actuators, Goldberg et al. (2019) presents a model for a caterpillar-inspired soft robot actuated by SMA, leveraging the work by (Bergou et al., 2008) related to the elastic rod theory and the Lagrange motion equation for systems of particles. A planar version of Bergou’s formulation is employed here to model the dynamics of a single SMA actuator. The actuator stiffness and damping parameters are first estimated and subsequently used to simulate the locomotion of the caterpillar robot fabricated from a series of SMA actuators. Interestingly, increasing the actuator stiffness by yields simulation displacements higher than the prototype by .
4.4.4 Bend & Stretch in Parallel Manipulators
The CRT also aided the design, statics modeling, and workspace analysis of an extensible rod-driven PCM with five DoFs at the tip (Wu and Shi, 2022a). A variant of the Levenberg-Marquardt method makes the PCM model suitable for real-time applications. Open-loop errors on the tip position in tracking tasks are lower than in PCC models. Also, they propose a numerical procedure for efficiently estimating partial workspace using a ROM.
@@ 4.5 Bend & Twist & Stretch
Five contributions explored bending, twisting, and stretching combined. For cable-driven actuation, it was presented a general geometrically exact static CRT for CSM (Renda and Laschi, 2012). The work highlights the role of torsion deformation. The variety of the design parameters allows the simulation of a wide range of continuum manipulators and the coupled tendon drive for multiple sections. Similarly, a cable-driven conic CSM made of silicone was modeled by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials (Wang et al., 2016). An experiment with the arm fixed at the base and oriented horizontally, subject to gravity, without actuation (), validates the viscoelastic model, which matches the experimental data better than a pure elastic model. Finally, the validation executed 2D and 3D motions actuating different cable combinations (). Also, a discrete Cosserat rod combined with a FEM described the dynamics of cable-driven soft robots (Adagolodjo et al., 2021). In particular, they use FEM to model the robot structure and Cosserat for the robot actuation. The implicit backward Euler time stepper integrated the differential equations for numerical stability. Direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods.
Aimed at colonoscopy, the CRT was employed to reconstruct the shape of a continuum endoscope given the force exerted on the distal end (Ryu et al., 2018). The experimental validation shows that longer endoscopes produce higher tip errors.
Also, Cosserat rods modeled a six-link PCM (Ghafoori and Khalaji, 2020). The model, initially validated for a single rod obtaining an average tip position error of , is applied to a six-link robot with good performance.
In a recent work (Roshanfar et al., 2023), the authors modelled and experimentally validated a soft robot with combined pneumatic and cable actuation. The robot is modelled using CRT where the tangent elastic modulus is expressed as a function of the internal pressure. In addition, the hyperelastic constitutive model was integrated into the Cosserat framework.
@@ 4.6 Bend & Stretch & Shear
Two works combine bending, stretching and shearing. For multi-backbone continuum robots, it was proposed a variable curvature statics-kinematic Cosserat rod model incorporating multi-backbone structural constraints (Chen et al., 2021). The model is validated on a robot with two multi-backbone continuum segments, each with four nitinol rods. The push-pull actuation bends the continuum robot. The experiments evaluate the tip positioning accuracy for circular trajectories and the shape discrepancy when the robot is subject to an end-point load in two configurations. They also implement the inverse kinematics in a real-time open-loop controller for a rectangular path following, with and without tip load. The proposed model outperforms the PCC.
Differently, a hybrid model combining the classical CRT with a data-driven stiffness estimation was developed (Bartholdt et al., 2021). Rather than solving numerically nonlinear constitutive equations, the model relies on a work point-dependent linear stress-strain relationship for computational efficiency. Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using artificial neural networks.
@@ 4.7 Bend & Twist & Shear
The only work that performed an experimental validation for the compound class of bending, torsion, and shearing is by Grazioso et al. (2019), who proposes a novel FEM-inspired spatial discretization technique. In particular, the rod is divided into several nodes, defining a helicoidal shape function for the interpolation. Thanks to the specific choice of the shape function, the Cosserat rod dynamics was written with Lie Groups. The method, implemented in a new simulator called SimSOFT, is validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions. In addition, they performed an experimental validation on the Princeton arm to benchmark the coupling of bending, torsion, and shear. The validation results report excellent performance in terms of accuracy and computational efficiency.
Bentley et al. (2023) employ a Cosserat string formulation to derive a tissue and needle force model. It describes normal and frictional forces along the shaft as a function of the planned needle path, friction model and parameters, and piercing force. The proposed force model and associated cost metric are safer and more clinically relevant for motion planning. They fit and validate the model through physical needle robot experiments in a gel phantom. The force model defines a bottleneck cost function for motion planning, which is evaluated against a path-length cost function in random environments, achieving a 62% peak force reduction.
@@ 4.8 Stretch & Shear
Combined stretching and shearing deformations are characteristic of peristaltic locomotion (Fig. 5). The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (Hemingway and O’Reilly, 2021). Incompressibility is enforced as an internal constraint using Green and Naghdi’s theory of a directed rod. Two linearly elastic isotropic material models are assumed. The material parameters are identified from experiments for small deformations. Motion effects from actuators or muscle contraction are modeled as external compressive loads using a doublet function for an assigned centerline force and a uniform pressure for a pair of assigned director forces. The method is showcased in a simulation of a soft robotic device.

### v2

@@ 6.3 Bend & Twist
| Buckling | sipos2020longest | Study relationship between buckling and on simulated cantilever rods. | CRT | |
| Bend & Stretch | Cable | xiao2023kinematics | Addressed stiffness regulation for a soft arm with a sliding backbone, achieving +57.7% stiffness. | CRT |
| chen2022model | Augments the PCC with a rod model to quickly solve the CSM deformation under . | EBRT | ||
| Pneumatic | sadati2017control | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. | CRT | |
| gilbert2019validation | Experimentally validated of a CSM model on bending motions generated by square pulse actuations. | CRT | ||
| Garbulinski2022bending | Characterized bending, natural frequency, and damping of a cantilevered extensile artificial muscle. | EBRT | ||
| alessi2023ablation | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | CRT | ||
| dou2023design | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | EBRT | ||
| Smart | goldberg2019planar | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. | CRT | |
| Rod-driven | wu2022design | Design, statics modeling, and numerical workspace analysis of an extensible rod-driven PCM. CRT | ||
| Bend & Twist & Stretch | Cable | renda2012general | Simulated the coupled tendon drive for multi-section CSM highlighting the role of twist. | CRT |
| wang2016three | Compared the Kelvin viscoelastic material model vs a purely elastic model on a real siliconic arm. | |||
| adagolodjo2021coupling | Combined FEM to model the robot body with CRT to model actuation. | |||
| Roshanfar2023Hyperelastic | Investigated pneumatic stiffness regulation on a cable-driven CSM using the hyperelastic material model. | |||
| Tubular | ryu2018application | Reconstructed the shape of a continuum endoscope given external forces exerted on the distal end . | CRT | |
| Parallel | ghafoori2020modeling | Modeled a six-link PCM starting from an experimental validation of a single elastic rod. | CRT | |
| Bend & Stretch & Shear | Tubular | chen2021variable | Experimentally validates a model for multi-backbone continuum robot. | CRT |
| Pneumatic | bartholdt2021parameter | Models a soft pneumatic actuator combining rods and ANN | ||
| Bend & Twist & Shear | — | grazioso2019geometrically | Proposed a novel FEM-inspired discretization technique and the SimSoft simulator. | CRT |
| Tubular | bentley2023safer | Proposed a force model and associated cost metric for safer and clinically relevant for motion planning. | ||
| Stretch & Shear | Bioinspired | hemingway2021continuous | Studied the peristaltic locomotion of the earthworm Lumbricus terrestris to model soft robots. | CRT |
| Stretch | Bioinspired | Hyousse2019distributed | Studied the peristaltic locomotion of a caterpillar-inspired soft robot. | CRT |
@@ 6.4 Bend & Stretch
Simultaneous bending and stretching behaviors are frequent in CSM actuated by cables or pneumatics (Figure 4). Notably, this category excludes twist and shear effects for several reasons: the actuators are linear and apply forces along the centerline, not generating torsional or shear stresses; interactions with the environment are either neglected or primarily induce axial and bending loads. As a result, the deformation behavior can be described through bending and axial stretching. We summarize several Bend & Stretch works in Table 2, whereas herein we focus the discussion on manufacturing errors, stiffness regulation, and fiber-reinforced actuators.
Manufacturing variability and material degradation are critical concerns in soft robotic systems, as they can significantly affect actuator behavior, model accuracy, and control performance. Variations in geometry, material stiffness, or the positioning of actuators during fabrication can cause substantial differences between the modeled behavior and the robot’s actual motion. While several experimentally validated rod models were developed for pneumatic soft arms (sadati2017geometry; sadati2017mechanics; sadati2017control; gilbert2019validation; Wang2022DynamicsMod; Lamping2023ANovel) and fingers (sachin2022analytical; flores2022soft), relatively few have focused specifically on the effects of manufacturing irregularities. Below, we discuss two works using Cosserat rod models to address manufacturing uncertainties in pneumatic soft robots. eugster2022soft described the kinematics of a soft arm using a nonlinear pressure-dependent constitutive law, the principle of virtual work, and modeled the actuator with strain energy functions. Manufacturing imperfections are considered by scaling the relations of extensional and bending stiffness and by a constant shift of the actuator position . Assuming the pressure dilates the chambers, the model also considers cross-section deformations with a pressure-dependent diameter . The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings, comparing constant or pressure-dependent chamber radius, with or without chamber repositioning, and with a linear or Ogden material law. Similarly, alessi2023ablation presented a dynamic model for a 3D-printed CSM, where the manufacturing variability caused the actuators to have different stretching capability when subject to equal actuation. They captured this effect by tuning the pressure-strain relation for each actuator introducing strain gains that tune the pressure-strain relation for each actuator
where is the interior actuator area. An ablation study conducted on dynamic motions investigated the contribution of different model components, uncovering that neglecting manufacturing uncertainties can degrade performance up to . In practice, manufacturing irregularities can affect every soft robotic system but remain unexplored outside the Bend & Stretch class. We attribute this to the fact that, for linear actuators, it is easier to quantify the expected motion and thus identify model discrepancies.
Stiffness modeling and regulation are central to advancing the adaptability and precision of soft robotic systems. However, in many designs, limitations such as single-mode actuation, the absence of antagonistic mechanisms, or the lack of a structural backbone can restrict the ability to modulate stiffness effectively. To address these challenges, Sun2018hybrid presented a novel design of a hybrid continuum robot whose actuators combine pneumatic muscles with embedded elastic rods. The robot changes its stiffness through a locking mechanism, switching between large-scale movement enabled by pneumatics and fine positioning enabled by push-pull of the rods. Stiffness tests revealed that the robot increases its structural stiffness by 65.1% during fine positioning in the locked mode and reduces the repetitive positioning error by 62.8%. To model this behavior, they improved a static KRT, calculating the elastic deformation through the minimal total potential energy principle in an optimal control framework. Similarly, xiao2023kinematics addressed stiffness regulation for a tendon-driven concentric tube soft robot with a Nitinol tube backbone, which can slide inside the soft robot body for pose or stiffness regulation. A Cosserat rod model was validated in different scenarios by varying the joint-space tendon input and task-space external contact force, achieving an average tip position error below 1% of the total length. Simulation studies demonstrate that the Nitinol backbone insertion can enhance the kinematic workspace and increase stiffness by 57.7%. Two manipulation case studies demonstrated the potential application of hybrid actuation for stiffness regulation. Another compliant rod-driven mechanism enabling stiffness regulation was introduced in a soft arm with pneumatic actuation (dou2023design). While an Euler-Bernoulli model was adapted to demonstrate the benefit of the mechanism on experiments of way-point tracking, the investigation of the stiffening behavior remains a future work.
Fiber-reinforced soft pneumatic actuators present significant modeling challenges due to nonlinear deformation under pressure. Several works leveraged variations of rod theories to more accurately capture their complex bending, elongation, and cross-sectional behaviors. sadati2017geometry; sadati2017mechanics; sadati2017control performed comprehensive investigations of several modeling approaches—all experimentally validated on the well-known STIFF-FLOP arm—addressing the challenge of cross-section deformation, captured analytically through geometry deformation and CRT. This approach outperformed constant curvature approaches by and variable curvature by . More recently, Berthold2022investigation introduced a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs. Specifically, they considered the compression effects of lateral pneumatic chamber walls by converting the inner chamber pressure into an equivalent force. Their model, validated against FEM simulations, obtained a mean tip position error up to . Conversely, Hanza2023mechanics accounted for the effect of radial pressure by proposing an inhomogeneous Cosserat rod model with a nonlinear strain-force relationship. The proposed method outperformed the standard CRT by about . Meanwhile, LiuShengkai2023Modeling explored the effects of self-gravity and external loads on the configuration of an actuator with a semicircular cross-section through an EBRT.
@@ 6.5 Bend & Twist & Stretch
A few works explored the combination of bending, twisting, and stretching, especially in cable-driven systems. renda2012general presented a general geometrically exact static CRT for cable-driven CSM. The work contributed with:
The simulation of coupled tendon drive for multiple sections, enabled by various design parameters.
Highlights of the role of torsion deformation.
Although the study was conducted in simulation only, it provided a solid foundation for future full dynamic analyses (Sec. 6.2). Similarly, wang2016three modeled a cable-driven conic CSM made of silicone by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials. The experiments with real-world data included:
Validation of the viscoelastic model with the arm fixed in a cantilever fashion, subject to gravity, without actuation ().
Execution of 2D and 3D motions actuating different cable combinations ().
They found that the Kelvin model matched the experimental data better than a pure elastic model. Also, adagolodjo2021coupling described the dynamics of cable-driven soft robots combining
FEM to model the robot structure, and
a discrete Cosserat rod to model the robot actuation.
The implicit backward Euler time stepper integrated the differential equations for numerical stability, while direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods. Recently, Roshanfar2023Hyperelastic investigated stiffness regulation on a cable-driven CSM equipped with a central linear pneumatic chamber. They captured robot motion and stiffening behavior by augmenting the classic Cosserat rod with the hyperelastic material model.
In conclusion, while the Bend & Twist & Stretch class comprises fewer works than the top four deformation classes, it has consistent contributions to cable-driven systems. This class could be considered a special case of the All Deformation, where the shearing strain was negligible and neglected. Other works in this class are reported in Table 2.
@@ 6.6 Bend & Stretch & Shear
The Bend & Stretch & Shear deformation class, similar to the previous class, emerges by neglecting torsion. chen2021variable proposed a variable curvature static-kinematic Cosserat rod model for multi-backbone continuum robots, incorporating multi-backbone structural constraints. The model is validated on a robot with two multi-backbone continuum segments, each with four nitinol rods. The push-pull actuation bends the continuum robot and transmits shear forces. The experiments evaluate the tip positioning accuracy for circular trajectories and the shape discrepancy when the robot is subject to an end-point load in two configurations. They also implement the inverse kinematics in a real-time open-loop controller for a rectangular path following, with and without tip load. The proposed model outperforms the PCC. Note that the torsion-free assumption did not hold in all experiments. Indeed, when out-of-plane external forces were applied to cause torsion, the tip position error raised to 19% of the length. Differently, bartholdt2021parameter developed a hybrid model combining the classical CRT with a data-driven stiffness estimation. Rather than solving numerically nonlinear constitutive equations, the model relies on a work point-dependent linear stress-strain relationship for computational efficiency. Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using ANN.
@@ 6.7 Bend & Twist & Shear
Another special case is when robots bend, twist, and shear, without stretching. In this context, grazioso2019geometrically proposed a novel FEM-inspired spatial discretization technique where the rod is divided into several nodes, defining a helicoidal shape function for the interpolation. Thanks to the specific choice of the shape function, they formulated the Cosserat rod dynamics with Lie Groups. The proposed method, also implemented in a new simulator called SimSOFT, was validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions. In addition, they experimentally validated the coupling between bending, torsion, and shear on the Princeton arm benchmark, achieving excellent accuracy and computational efficiency. bentley2023safer employ a Cosserat string formulation to derive a tissue and needle force model. It describes normal and frictional forces along the shaft as a function of the planned needle path, friction model and parameters, and piercing force. The proposed force model and associated cost metric are safer and more clinically relevant for motion planning. They fit and validate the model through physical needle robot experiments in a gel phantom. The force model defines a bottleneck cost function for motion planning, which is evaluated against a path-length cost function in random environments, achieving a 62% peak force reduction.
@@ 6.8 Stretch & Shear
Combined stretching and shearing deformations are characteristic of peristaltic locomotion. The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (hemingway2021continuous). Incompressibility is enforced as an internal constraint using Green and Naghdi’s theory of a directed rod. Two linearly elastic isotropic material models are assumed. The material parameters are identified from experiments for small deformations. Motion effects from actuators or muscle contraction are modeled as external compressive loads using a doublet function for an assigned centerline force and a uniform pressure for a pair of assigned director forces. The method is showcased in a simulation of a soft robotic device.